# Gaussian On-the-Fly Splatting: A Progressive Framework for Robust Near Real-Time 3DGS Optimization

### v1

@@ IV-D Adaptively Learning Rate Updating
The progressive training in On-the-Fly GS may result in inconsistency in local quality within the 3DGS field. To address this issue, we dynamically update the learning rate based on the rendering quality of each image, instead of updating the learning rate based on the overall quality of the 3DGS field (this was applied in the original 3DGS). Specifically, for image $j$, if it has been trained for $\text{IT}_{j}$ iterations, with the initial learning rate ILr and the final learning rate FLr, the current learning rate $\text{LogLr}_{j}$ for the image $j$ can be calculated as follows:
$\text{LogLr}_{j}=e^{\ln\left(\text{ILr}\right)\times\left(1-t_{i}\right)+\ln\left(\text{FLr}\right)\times t_{i}},$ (8)8( 8 )
$t_{i}=\left\{\begin{matrix}1,&\text{IT}_{j}>\text{Mean}\\ \frac{\text{IT}_{j}}{\text{Mean}},&\text{IT}_{j}\leq\text{Mean}\end{matrix}\right.,$ (9)9( 9 )
where Mean is a predefined parameter that represents the number of training iterations required for an image to achieve satisfactory rendering quality. However, determining the learning rate based on the already trained iterations of a single image is not rigorous. This is because other images overlapping with image $j$ might already achieve satisfactory rendering quality. Thus, when determining the current learning rate, it makes sense to also consider the overlapping images among the registered images. If the set of overlapping images $O_{I}$ for the current image $j$ contains $RN$ images (i.e., for any $i\in O_{I},M_{i,j}>0$). The learning rate $\text{Lr}_{j}$ for image $j$ can be calculated as follows:
$\text{Lr}_{j}=\frac{\text{LogLr}_{j}+{\textstyle\sum_{i=1}^{RN}}M_{i,j}\times\text{LogLr}_{i}}{RN+1}.$ (10)10( 10 )
The proposed learning rate updating method considers the rendering quality of the 3DGS field within the local scene. If the local scene corresponding to the current image $j$ has achieved high rendering quality, the learning rate $\text{Lr}_{j}$ is supposed to be lower. This is because multiple images overlapping with image $j$ have been trained for more iterations (even if image $j$ is newly added), as shown in Fig. 4 (left). If the local scene corresponding to the image $j$ performs poor rendering quality, which indicates that both image $j$ and its neighboring connected images are only trained for fewer iterations, the learning rate $\text{Lr}_{j}$ should be higher, as shown in Fig. 4 (right).
@@ IV-E Load Balancing and 3DGS Runtime Optimization
Some local areas may undergo excessive Gaussian accumulation due to being visible across multiple images, while newly introduced scenes may suffer from insufficient Gaussian kernels. To address this uneven distribution of Gaussians, inspired by Adr-Gaussian[5], we employ a load balancing training. For the current training image $i$, we record the number of Gaussians $\text{GN}_{p}$ associated with each pixel $p$ in image $i$. Using the standard deviation of Gaussian distribution within the visible range of the image, we define a Load-Balancing-Loss $L_{\text{load}}$ as follows:
$L_{\text{load}}=\text{std}\left(\text{GN}_{p}\right),p\in i.$ (11)11( 11 )
@@
$\text{LOSS}=\lambda_{1}L_{1}+\lambda_{\text{ssim}}L_{\text{ssim}}+\lambda_{\text{load}}L_{\text{load}}.$ (12)12( 12 )
Subsequent experiments demonstrate that the load balancing strategy not only prevents the abnormal accumulation of Gaussians in some specific regions but also effectively reduces the total number of Gaussians in the 3DGGS field, thereby accelerating the overall training process. Furthermore, it facilitates the densification of Gaussians in newly added regions, promoting faster 3DGS training.
To further accelerate the training of the 3DGS field, we integrate the strategy proposed by Taming 3DGS[9]. Taming 3DGS identified that the per-pixel gradient accumulation method used in the original 3DGS leads to multiple threads contending for accessing the same locations and thus serialized atomic operations. To address this, we implemented the corresponding Splatting-based parallelism in the gradient accumulation method. In this implementation, gradients are accumulated directly for each Gaussian, significantly reducing training time by eliminating contention and enhancing parallel processing efficiency.
@@ V EXPERIMENT
TABLE I: Comparison with offline solutions at same rendering performance as on-the-fly GS. Best and second-best are highlighted in bold and underline.
Dataset ImageNum Methods Video Memory (MB) PSNR
COLMAP/
On-the-Fly SfM (min)
3DGS Training (min) Total Time (min)
Counter 240 3DGS 218.07 27.68 16.92 4.91 21.83
Taming-3DGS 221.38 1.38 18.20
Adr-Gaussian 121.56 3.76 20.68
Ours-Offline GS 119.06 3.14 20.06
On-the-Fly GS 118.36 14.85 \ 14.85+0.48
Train 301 3DGS 128.15 21.01 16.39 2.82 19.21
Taming-3DGS 135.44 1.31 17.70
Adr-Gaussian 69.73 2.18 18.57
Ours-Offline GS 75.61 1.44 17.83
On-the-Fly GS 159.68 19.21 \ 19.21+0.36
YD 125 3DGS 2124.37 22.04 12.55 13.58 26.13
Taming-3DGS 2183.01 8.31 20.85
Adr-Gaussian 1384.139 15.27 27.82
Ours-Offline GS 1084.87 6.48 19.03
On-the-Fly GS 699.34 7.70 \ 7.70+0.55
Palace 210 3DGS 100.95 21.41 14.17 10.49 24.66
Taming-3DGS 118.36 4.02 18.19
Adr-Gaussian 48.68 9.65 23.82
Ours-Offline GS 62.01 3.46 17.63
On-the-Fly GS 173.94 16.54 \ 16.54+0.38
@@ V-A Experimental Settings
Datasets and Metrics. To validate the efficacy of our On-the-Fly GS, we utilizes several datasets from the 3DGS[3], Mip-NeRF 360[37], On-the-Fly SfM[36, 38], and TUM-RGBD[42]. This contains three indoor scenes, two small-scale outdoor scenes, and one UAV image, which are dynamically processed as simulated sequential inputs according to the original stored order. Based on these datasets, the corresponding performance is evaluated from three aspects: 1) Training efficiency. Since On-the-Fly GS is capable of simultaneous estimation of image poses and 3DGS training, we record the time consumption of different methods starting from the moment image pose estimation begins until the final 3DGS field is obtained. 2) Rendering quality. We use the same common metrics as those used in 3DGS, including PNSR, SSIM and LPIPS. 3) Video Memory. We analyse the trained 3DGS field memory consumption of each method. In addition, for On-the-Fly GS, we record the time of the complete training process and evaluate the rendering quality. To compare with other offline-training methods, we first investigate the training time required to achieve the same rendering quality as our On-the-Fly GS, followed by exploring their default training iteration settings. All experiments are conducted under the same hardware configuration (Single NVIDIA RTX 4090 GPU).
Implementation details. The optimization process of our On-the-Fly GS includes three stages (see Fig. 2). In phase 1 of initialization, after On-the-Fly SfM completes pose estimation and sparse point cloud generation for the first $N_{0}=16$ images, we perform 2,000 training iterations on these images. In phase 2, for each newly acquired image, we first run 100 iterations of local optimization, followed by 100 iterations of global optimization. Additionally, the preset parameter Mean is 195, ILr is $41.6e-$ 4 and FLr is $61.6e-$ 6. In the last phase 3, an additional 1,000 iterations of fast global optimization are performed to further improve the overall model quality. The weight of Load-Balancing-Loss $\lambda_{\text{load}}$, L1-Loss $\lambda_{1}$ and SSIM-Loss $\lambda_{\text{ssim}}$ are set to 0.41, 0.47 and 0.12, respectively.
@@ V-B Comparisons with other SOTA methods
We compare the proposed On-the-Fly GS with four offline-training solutions (3DGS[3], Adr-Gaussian[5], Taming-3DGS[9] and Ours-Offline GS training). Ours-Offline GS refers to an offline training pipeline that incorporates Load Balancing and Splatting-based parallelism in backpropagation strategies, while keeping all other settings identical to 3DGS. COLMAP is employed for pose estimation and sparse point cloud generation for the offline-training methods, while On-the-Fly SfM is employed to estimate pose and generate sparse point cloud for our On-the-Fly GS.
Table I reports the corresponding metrics when the four offline training solutions achieve the same rendering performance as that can be generated by our On-the-Fly GS. It can be seen that On-the-Fly GS shows a significant advantage in terms of cost time. Apart from the time (before the +) required for On-the-Fly SfM to do online pose estimation and sparse point cloud generation, only a small amount of extra overload (after the +) is needed to complete the 3DGS training. Generally, it takes around 2-3 seconds for a newly fly-in image to finish the 3DGS optimization, which is sufficient for near real-time performance and demonstrates the capability of simultaneous image acquisition, pose estimation, and 3DGS training. On the other hand, the offline training solutions need to finish the Colmap processing before training. They also typically take longer for both pose estimation and 3DGS training, further leading to more total time for offline training. Furthermore, the model obtained through On-the-Fly GS has the smallest model size in some scenarios, which means that the proposed method is promising to enable lightweight 3DGS training.
To further investigate the discrepancy between our work and these offline solutions with default setting, the number of iterations for these offline methods is set to 30,000. As shown in Tab. II and Fig. 5, in general, the overall rendering quality of our On-the-Fly GS is around 80%–90% of that generated by these offline solutions. Nevertheless, our On-the-Fly GS demonstrates a significant advantage in terms of time efficiency, while also producing superior results in certain rendered images. Furthermore, as YD is a UAV image dataset, the 3DGS and Taming-3DGS can only run 11,000 and 10,500 iterations due to excessive memory consumption (the results presented in Tab. II and Fig. 5 correspond to this specific training setting), which also prove the effectiveness of our method in handling model size.
On-the-fly SfM[36] can also process sequential video frames, thus, we further tested two SLAM benchmarks and compared with several real-time NVS methods based on SLAM, including NICE-SLAM[39], Vox-Fusion[40], GS-SLAM[21], and MonoGS++[41]. Quantitative comparisons are shown in Table III. On-the-Fly GS typically performs the best or second best on the evaluation metrics.
@@
Figure 5: Rendering Results of On-the-Fly GS and Offline solutions with default settings
TABLE II: Comparison with offline solutions of default setting. Best and second-best are highlighted in bold and underline.
Methods Total Time (min) PSNR
Counter Train YD Palace Counter Train YD Palace
3DGS 37.35 26.06 25.03 39.16 30.37 24.64 22.09 23.98
Taming-3DGS 24.66 21.39 20.85 23.89 30.25 24.31 22.15 23.71
Adr-Gaussian 28.68 21.71 45.19 28.09 29.57 23.07 23.57 22.19
Ours-Offline GS 24.16 19.98 34.71 21.93 29.75 22.79 24.13 23.83
On-the-Fly GS 15.33 19.57 8.25 16.92 27.68 21.01 22.04 21.41
@@
TABLE III: Comparison with SLAM-based method. Best and second-best are highlighted in bold and underline.
Methods fr1_room fr1_desk
PSNR SSIM LPIPS PSNR SSIM LPIPS
NICE-SLAM 11.39 0.37 0.62 13.83 0.56 0.48
Vox-Fusion 14.20 0.56 0.55 15.79 0.64 0.52
MonoGS 14.76 0.52 0.60 17.31 0.65 0.38
MonoGS++ 22.16 0.77 0.31 22.85 0.82 0.60
Ours 22.17 0.84 0.32 24.23 0.86 0.26

### v2

@@ IV-D Adaptively Learning Rate Updating
The progressive training in On-the-Fly GS may result in inconsistency in local quality within the 3DGS field. To address this issue, we dynamically update the learning rate based on the rendering quality of each image, instead of updating the learning rate based on the overall quality of the 3DGS field (this was applied in the original 3DGS). Specifically, for image $j$, if it has been trained for $\text{T}_{j}$ iterations, with the initial learning rate $\eta_{j}^{0}$ and the final learning rate $\eta_{j}^{f}$, the current learning rate $\eta_{j}^{c}$ can be calculated as follows:
$\eta_{j}^{c}=e^{\ln\left(\eta_{j}^{0}\right)\times\left(1-t_{i}\right)+\ln\left(\eta_{j}^{f}\right)\times t_{i}},$ (6)( 6 )
$t_{i}=\left\{\begin{matrix}1,&\text{T}_{j}>\text{T}_{a}\\ \frac{\text{T}_{j}}{\text{T}_{a}},&\text{T}_{j}\leq\text{T}_{a}\end{matrix}\right.,$ (7)( 7 )
where $\text{T}_{a}$ is a predefined parameter that represents the number of training iterations required for an image to achieve satisfactory rendering quality. However, determining the learning rate based on the already trained iterations of a single image is not rigorous. This is because other images overlapping with image $j$ might already achieve satisfactory rendering quality. Thus, when determining the current learning rate, it makes sense to also consider the overlapping images among the registered images. If the set of overlapping images $\mathcal{I}$ for the current image $j$ contains $N_{\mathcal{I}}$ images (i.e., for any $i\in\mathcal{I},M_{i,j}>0$). The learning rate $\eta_{j}^{\mathcal{I}}$ for image $j$ can be calculated as follows:
$\eta_{j}^{\mathcal{I}}=\frac{\eta_{j}^{c}+{\textstyle\sum_{i=1}^{N_{\mathcal{I}}}}M_{i,j}\times\eta_{i}^{c}}{N_{\mathcal{I}}+1}.$ (8)( 8 )
The proposed learning rate updating method considers the rendering quality of the 3DGS field within the local scene. If the local scene corresponding to the current image $j$ has achieved high rendering quality, the learning rate $\eta_{j}^{\mathcal{I}}$ is supposed to be lower. This is because multiple images overlapping with image $j$ have been trained for more iterations (even if image $j$ is newly added), as shown in Fig. 4 (left). If the local scene corresponding to the image $j$ performs poor rendering quality, which indicates that both image $j$ and its neighboring connected images are only trained for fewer iterations, the learning rate $\eta_{j}^{\mathcal{I}}$ should be higher, as shown in Fig. 4 (right).
@@ IV-E Load Balancing and 3DGS Runtime Optimization
Some local areas may undergo excessive Gaussian accumulation due to being visible across multiple images, while newly introduced scenes may suffer from insufficient Gaussian kernels. To address this uneven distribution of Gaussians, inspired by Adr-Gaussian[5], we employ a load-balancing training. For the current training image $i$, we record the number of Gaussians $\text{G}_{p}$ associated with each pixel $p$ in image $i$. Using the standard deviation of the Gaussian distribution within the visible range of the image, we define a Load-Balancing-Loss $L_{\text{load}}$ as follows:
$L_{\text{load}}=\text{std}\left(\text{G}_{p}\right),p\in i.$ (9)( 9 )
@@
$\text{L}_{\text{total}}=\lambda_{1}L_{1}+\lambda_{\text{ssim}}L_{\text{ssim}}+\lambda_{\text{load}}L_{\text{load}}.$ (10)( 10 )
Subsequent experiments demonstrate that the load-balancing strategy not only prevents the abnormal accumulation of Gaussians in some specific regions but also effectively reduces the total number of Gaussians in the 3DGS field, thereby accelerating the overall training process. Furthermore, it facilitates the densification of Gaussians in newly added regions, promoting faster 3DGS training.
To further accelerate the training of the 3DGS field, we integrate the strategy proposed by Taming 3DGS[7]. Taming 3DGS identifies that the per-pixel gradient accumulation method used in the original 3DGS leads to multiple threads contending for accessing the same locations and thus serialized atomic operations. To address this, we implemented the corresponding Splatting-based parallelism in the gradient accumulation method. In this implementation, gradients are accumulated directly for each Gaussian, significantly reducing training time by eliminating contention and enhancing parallel processing efficiency.
@@ V-A Experimental Settings
Datasets and Metrics. To validate the efficacy of our On-the-Fly GS, we utilizes several datasets from the 3DGS (RGB)[3], Mip-NeRF 360 (RGB)[33], On-the-Fly SfM (RGB)[22, 1], and Replica (RGB-D)[34]. These contain three indoor scenes, three small-scale outdoor scenes, three UAV datasets and eight VSLAM datasets, which are dynamically processed as simulated sequential inputs according to the original stored order. Based on these datasets, the corresponding performance is evaluated from three aspects: Training efficiency, Rendering quality, and Model size. All experiments are conducted under the same hardware configuration (Single NVIDIA RTX 4090 GPU).
TABLE I: Comparison on RGB datasets. Best and second-best are highlighted in bold and underlined.
Type Indoor Outdoor UAV
Datasets skfx kitchen counter train Temple JYYL WHU npu phantom3 Avg.
Image Number 60 279 240 301 302 390 242 300 399
Pose Estimation (s) 95.48 1037.27 615.67 813.98 1068.51 1373.87 820.06 808.21 1183.37
Adr-Gaussian
PSNR (dB) 31.07 28.45 26.81 19.94 18.68 19.85 24.99 26.85 26.83 24.83
SSIM 0.87 0.89 0.86 0.69 0.67 0.69 0.82 0.81 0.83 0.79
LPIPS 0.28 0.15 0.24 0.36 0.57 0.47 0.17 0.29 0.26 0.31
Model Size (MB) 323.4 192.5 104.5 63.2 21.3 92.8 839.2 895.5 809.1 371.3
Ad-Opt Time (s) 176.36 245.96 207.96 109.36 208.59 280.58 268.09 268.98 279.15 227.23
Taming-3DGS
PSNR (dB) 32.43 29.41 27.52 20.57 19.39 20.16 25.33 27.25 27.16 25.47
SSIM 0.89 0.90 0.88 0.72 0.69 0.71 0.84 0.83 0.84 0.81
LPIPS 0.22 0.13 0.21 0.33 0.54 0.44 0.14 0.26 0.23 0.28
Model Size (MB) 652.7 386.9 217.6 149.7 63.1 241.8 1648.4 1784.8 1596.7 749.1
Ad-Opt Time (s) 113.65 155.67 123.87 75.34 88.39 103.47 245.62 236.58 240.92 153.72
Taming-3DGS
PSNR (dB) 37.84 33.01 30.16 25.35 22.58 23.93 \ \ \ 28.81
SSIM 0.95 0.94 0.92 0.86 0.77 0.82 \ \ \ 0.88
LPIPS 0.09 0.08 0.15 0.15 0.37 0.25 \ \ \ 0.18
Model Size (MB) 950.4 417.7 250.2 233.1 122.6 504.8 \ \ \ 413.1
Ad-opt Time (s) 950.34 953.09 475.13 350.17 343.09 612.72 \ \ \ 614.09
Ours PSNR (dB) 31.93 30.39 28.97 22.82 21.72 22.71 27.93 33.29 31.46 27.91
SSIM 0.86 0.91 0.89 0.80 0.75 0.79 0.87 0.93 0.92 0.86
LPIPS 0.29 0.11 0.19 0.21 0.39 0.28 0.11 0.11 0.12 0.20
Model Size (MB) 113.1 176.4 115.9 137.4 133.1 416.9 534.4 672.3 831.9 347.9
Ad-Opt Time (s) 53.74 66.06 46.45 46.67 58.24 76.81 91.61 105.97 114.55 73.34
FPS 0.63 0.27 0.39 0.37 0.31 0.28 0.30 0.37 0.34 0.36
GS-SLAM
PSNR (dB) \ 17.59 18.19 16.34 \ 18.25 20.01 21.48 20.57 18.92
SSIM \ 0.48 0.63 0.49 \ 0.66 0.48 0.65 0.66 0.58
LPIPS \ 0.76 0.74 0.72 \ 0.62 0.58 0.68 0.75 0.69
Model Size (MB) \ 43.6 22.6 23.7 \ 26.1 44.6 38.2 29.9 32.67
FPS \ 0.19 0.19 0.37 \ 0.26 0.23 0.21 0.26 0.24
Finish Tracking \
[…]
@@ V-B Comparisons with other SOTA methods
We compare On-the-Fly GS with two efficient offline-training solutions (Adr-Gaussian[5] and Taming-3DGS[7]) and two popular SLAM-based 3DGS methods (GS-SLAM[19] and RTG-SLAM[18]). Since RTG-SLAM only accepts RGB-D input, comparisons on RGB datasets are conducted solely among the offline training solutions and GS-SLAM. RTG-SLAM is included in the comparison only when using RGB-D datasets. Unless otherwise specified, all offline training solutions and SLAM-based solutions are evaluated using their default parameter settings. ‘
@@
Figure 5: Rendering Results of On-the-Fly GS and other solutions with default settings
Results on RGB Datasets. To conduct a fair evaluation, the offline-training solutions also use the SfM results from On-the-Fly SfM, but after all images are solved. The time consumption for online SfM is reported in Table I. As On-the-Fly GS and the offline-training solutions undergo the same pose estimation process, we only record the cost time from the completion of pose estimation to the completion of 3DGS field optimization (Ad-Opt Time). Since On-the-Fly GS performs training concurrently with the online SfM, in fact, only the cost time for the additional optimization in Phase 3 is recorded. In contrast, the offline solutions report their total execution time. For GS-SLAM, we report its FPS using the default setting. In our On-the-Fly GS, the reported FPS takes the entire duration of both online SfM and 3DGS field optimization into account. Moreover, since the degree of overlap between image pairs in the RGB datasets is significantly lower than that of typical SLAM datasets, each image is regarded as a keyframe when training GS-SLAM. Table I and Fig. 5 show the comparison results on RGB datasets.
@@
TABLE II: Comparison on RGB-D datasets. Best and second-best are highlighted in bold and underlined.
Method Modality Metric room0 room1 room2 office0 office1 office2 office3 office4 Avg.
GS-SLAM RGB PSNR (dB) 29.31 27.58 29.24 33.88 35.51 25.24 28.79 27.89 29.89
SSIM 0.89 0.84 0.89 0.93 0.94 0.88 0.89 0.91 0.89
LPIPS 0.24 0.32 0.23 0.18 0.16 0.28 0.18 0.24 0.22
FPS 0.89 0.92 0.93 0.95 1.05 0.89 0.93 0.91 0.93
ATE-MSE (m) 0.0849 0.4468 0.1897 0.2014 0.2144 0.5454 0.1664 0.8466 0.1869
RTG-SLAM RGB-D PSNR (dB) 26.45 30.04 31.31 36.82 37.37 30.05 30.18 33.28 31.94
SSIM 0.92 0.94 0.95 0.98 0.99 0.96 0.97 0.97 0.96
LPIPS 0.28 0.22 0.19 0.10 0.11 0.19 0.18 0.16 0.18
FPS 2.71 2.94 2.73 2.71 3.05 3.13 2.94 2.94 2.89
ATE-MSE (m) 0.0019 0.0019 0.0011 0.0015 0.0013 0.0024 0.0023 0.0025 0.0019
Ours RGB PSNR (dB) 31.79 33.67 34.65 39.44 38.77 34.03 35.38 36.65 35.55
SSIM 0.94 0.95 0.96 0.97 0.96 0.95 0.96 0.96 0.96
LPIPS 0.12 0.11 0.11 0.09 0.14 0.13 0.09 0.10 0.11
FPS 2.01 2.23 1.99 3.25 3.56 1.86 1.88 2.12 2.36
ATE-MSE (m) 0.0035 0.0054 0.0025 0.0057 0.0053 0.0070 0.0024 0.0075 0.0049
Compared to the offline solutions with 7K iterations, our On-the-Fly GS achieves significantly higher rendering quality as well as explicitly less cost time. This indicates that, thanks to the proposed Local & Semi-Global training strategy, our On-the-Fly GS can quickly achieve high-quality rendering with just a small number of additional iterations. Furthermore, On-the-Fly GS exhibits a pronounced advantage on the UAV dataset, significantly outperforming other solutions in both training efficiency and rendering quality. These results underscore the strong potential of the proposed approach for large-scale UAV image dataset applications. With 30K training iterations, Taming-3DGS is unable to complete training on the UAV image datasets due to excessive memory consumption. As for other scenes, On-the-Fly GS consistently achieves over 80% of the rendering quality delivered by Taming-3DGS, while exhibiting clear advantages in terms of model size and cost time.
Due to the significantly low overlapping degree between consecutive image pairs in the datasets presented in Table I, GS-SLAM exhibits extremely poor camera tracking performance across all the scenes, and even fails to track on some datasets (skfx and Temple). Additionally, the tested RGB datasets have high resolutions and very complex scene content, which leads to an increase in the training time for GS-SLAM and a substantially lower rendering quality compared to other methods. These results reveal the potential limitations of SLAM-based 3DGS methods on handling datasets without a reliable depth map or containing consecutive image pairs of low overlapping degree.
Results on RGB-D Datasets. To compare the performance of SLAM-based 3DGS method, the Replica dataset with depth priors is tested on the two popular pipelines - GS-SLAM[19] and RTG-SLAM[18]. Since these two methods’ solutions perform 3DGS field optimization using only the selected keyframes, we retain the selected keyframes from GS-SLAM, which are subsequently used as input for our On-the-Fly GS.
Table II shows the comparison results on the RGB-D dataset. It can be seen that, despite only RGB information from keyframes is used as input, the proposed On-the-Fly GS achieves the best or the second-best rendering quality among all evaluated methods and datasets. Although On-the-Fly GS is slightly inferior to RTG-SLAM in terms of time efficiency and tracking accuracy, our method attains superior rendering quality without requiring depth and does not rely on sequential video frames as input. Instead, it only requires that newly added images have a certain degree of overlap with any previously registered views. These characteristics make On-the-Fly GS more broadly applicable than existing SLAM-based solutions.